### v1

Abstract
Urdu, as a low-resource language, lacks effective semantic content recommendation systems, particularly in the domain of personalized news retrieval. Existing approaches largely rely on lexical matching or language-agnostic techniques, which struggle to capture semantic intent and perform poorly under varying query lengths and information needs. This limitation results in reduced relevance and adaptability in Urdu content recommendation. We propose ULTRA (Urdu Language Transformer-based Recommendation Architecture), an adaptive semantic recommendation framework designed to address these challenges. ULTRA introduces a dual-embedding architecture with a query-length aware routing mechanism that dynamically distinguishes between short, intent-focused queries and longer, context-rich queries. Based on a threshold-driven decision process, user queries are routed to specialized semantic pipelines optimized for either title/headline-level or full-content/document level representations, ensuring appropriate semantic granularity during retrieval. The proposed system leverages transformer-based embeddings and optimized pooling strategies to move beyond surface-level keyword matching and enable context-aware similarity search. Extensive experiments conducted on a large-scale Urdu news corpus demonstrate that the proposed architecture consistently improves recommendation relevance across diverse query types. Results show gains in precision above 90% compared to single-pipeline baselines, highlighting the effectiveness of query-adaptive semantic alignment for low-resource languages. The findings establish ULTRA as a robust and generalizable content recommendation architecture, offering practical design insights for semantic retrieval systems in low-resource language settings.
Index Terms:
Urdu Content Recommendation System, Intelligent Information Retrieval, Semantic Search, Urdu NLP, Dual-Embedding Framework, Query Length Optimization,
[…]
=-15pt
I Introduction
Text-based content recommendation systems have become indispensable in the digital era, transforming how users discover and consume information amidst an overwhelming volume of content [1, 2]. By leveraging user behavior, interests, and search queries, these systems curate personalized content feeds, enhancing engagement and information accessibility. At their core, modern systems rely on sophisticated semantic understanding to transcend simple keyword matching, aiming to grasp user intent and the contextual meaning of content for accurate alignment [3, 4]. Figure 2 illustrates the evolution of text-based recommendation and retrieval techniques from traditional lexical methods to modern semantic approaches.
Figure 1: Evolution of text-based recommendation & retrieval techniques
This challenge of semantic recommendation is not uniform; it is heavily modulated by linguistic and computational resource disparities [5]. For high-resource languages like English, advanced transformer-based architectures have achieved remarkable success [6, 7, 8]. However, for linguistically rich but computationally low-resource languages such as Urdu, the landscape is starkly different. Urdu, with large number of speakers worldwide and written in the complex Nastaliq script, presents unique morphological and syntactic challenges [9, 10]. Its context sensitivity, diglossia, and significant code-mixing, particularly with English and Roman Urdu in digital spaces, complicate automated semantic analysis [11, 12, 13]. This creates a critical barrier to accessing relevant digital content for a vast population.
Existing research aimed at Urdu text-based content recommendation has primarily explored foundational techniques. Early systems and some contemporary approaches rely on traditional lexical methods, such as TF-IDF with cosine similarity, which are fundamentally limited by their dependence on surface-level keyword overlap [14, 15]. Acknowledging this limitation, subsequent work has integrated pre-trained transformer models like BERT to generate semantic embeddings, marking a significant step forward [16]. Further innovations have focused on enriching content representation through entity-based techniques, such as wikification and knowledge graphs, aiming to connect content items via named entities and their relationships [17, 18]. While these methods improve upon purely lexical systems, they exhibit notable shortcomings that hinder optimal performance in real-world scenarios [24, 25, 26].
A principal drawback of these existing approaches is their rigid, one-size-fits-all architecture. They typically process all user queries through a single, uniform matching pipeline, whether against headlines, full articles, or entity graphs [19]. This design fails to account for the inherent diversity in user query patterns, which range from short, intent-focused searches (e.g., ”result of elections”) as in Figure 2,
Figure 2: Example of a short, intent-focused Urdu query
to long, descriptive inquiries seeking specific contextual details (e.g.,“regarding the results of Pakistan’s 2024 national elections, which political parties secured how many seats in different provinces and constituencies?”) as shown in Figure 3. [20].
Figure 3: Example of a long, context-rich Urdu query
However, remember long queries are not limited to user-entered inputs alone; a detailed definition of long queries, as established in our work, is provided in Section II.
Consequently, a concise query may be inadequately matched against verbose article bodies, losing precision, while a detailed query may not fully leverage the rich semantics embedded in full-text content, compromising recall [21]. Furthermore, scalability remains a pervasive concern, as the computational cost of performing real-time similarity searches on high-dimensional transformer embeddings across large and growing content corpora is often prohibitive, limiting practical deployment [22, 23]. There is a clear, unaddressed need for an adaptive system that dynamically tailors its retrieval strategy to the nature of the user’s query [27, 28, 29].
To address these critical limitations, we propose ULTRA (Urdu Language Transformer Based Recommendation Architecture), a novel dual-pathway framework designed for adaptive, high-precision semantic text-based content recommendation in Urdu. The core innovation of ULTRA is an intelligent routing mechanism that dynamically directs user queries to specialized semantic pipelines based on query characteristics. This ensures the granularity of the user’s informational need is aligned with the most appropriate level of document representation. By leveraging state-of-the-art transformer models fine-tuned for Urdu within an optimized vector search infrastructure, ULTRA achieves genuine semantic comprehension beyond lexical coincidence, while maintaining efficiency for large-scale deployment [29].
II Our Novel Contributions Explained through Use Cases
Figure 4 provides an abstract representation of the ULTRA architecture, depicting the structural organization and functional relationships between its core components. Building upon this architectural foundation, the main contributions of this work can be summarized as follows:
II-A Use Cases in Text Based Content Recommendation Systems
When recommending text-based content on content-based recommendation systems such as news websites, e-commerce sites, blog platforms, and digital libraries, two primary use cases emerge: (1) users enter a query into a search bar, triggering recommendations aligned with the query; and (2) users are engaged with a specific webpage, such as reading a news article, blog post, or product description on an e-commerce site, where related content is suggested below the current page.
Our proposed architecture accommodates both scenarios. In the case of a user entering a query via the search bar—typically consisting of keywords and thus classified as a short query—the input is routed to a pipeline optimized for short queries. Conversely, when recommending content based on the webpage the user is currently viewing, the text on that specific web page serves as the query, which falls under the long query category. Recommendations are then generated relative to the topic the user is actively engaged with. However users can also type long queries in search bar; such inputs would similarly be directed to the long-query-optimized pipeline. To distinguish between short and long queries, we establish a predefined threshold value. The methodology for selecting and determining this threshold is detailed in section Section IV-A2.
II-B Dual Pipeline Approach
The distinction between short and long queries is motivated by experimental findings detailed in Section V. Our evaluation revealed that short queries, composed of a few keywords, exhibit poor precision when matched against embeddings of lengthy content (e.g., complete articles or blog posts). However, matching these queries against headline embeddings, which are comparable in length and linguistic structure, yielded substantially higher precision.
These observations drove our dual-pipeline architectural design. Rather than matching compact queries directly against full document embeddings, we propose an alternative strategy: (1) encode content semantics via headline embeddings, (2) retrieve relevant headlines based on query-headline similarity, and (3) recommend the complete documents associated with top-ranked headlines. This approach exploits the semantic and structural correspondence between query brevity and headline conciseness, thereby improving recommendation accuracy for Urdu text corpora.
II-C Major Points
  • We propose ULTRA, a novel adaptive recommendation architecture that introduces a query-length threshold-based routing mechanism, as established in Section II-A and II-B, to dynamically optimize semantic matching, a novel approach for Urdu text-based content recommendation.
  • We conduct an extensive empirical evaluation of embedding generation and pooling strategies, establishing optimal configurations for representing both title-level and full document-level Urdu content.
  • We implement and rigorously compare multiple dimensionality reduction techniques as well as reduced dimension sizes, identifying the most effective method and dimension size for compressing semantic embeddings to enable efficient, large-scale similarity search without significant fidelity loss.
  • We demonstrate through comprehensive experiments that ULTRA sets a new performance benchmark, consistently achieving precision rates exceeding 90% and significantly outperforming existing non-adaptive semantic and lexical baselines for Urdu text based content recommender systems.
Figure 4: Architecture diagram of proposed ULTRA recommendation framework
III Related Work
Text-based content recommendation systems have evolved significantly, leveraging advancements in natural language processing (NLP) to provide personalized experiences across various domains, including news, articles, social media, and e-commerce. Early systems primarily focused on content-based filtering, where recommendations are generated based on the similarity between item features and user preferences, often using vector space models [1]. For instance, a foundational survey on recommender systems highlighted the transition from simple keyword-based methods to more sophisticated hybrid approaches that incorporate user behavior and contextual data [30]. These systems aim to address information overload by curating relevant text content, but they often struggle with semantic depth, especially in diverse linguistic contexts [2].
In high-resource languages like English, content recommendation has benefited from large-scale datasets and advanced models. A comprehensive review of text-based recommendation systems emphasized the role of feature selection techniques, such as word embeddings, in improving accuracy over traditional methods [14]. This survey analyzed over 100 studies, concluding that hybrid models combining content-based and collaborative filtering yield better personalization [31]. Another key work proposed a content-based book recommender using machine learning for text categorization, demonstrating how information extraction can enhance relevance in textual domains [32]. However, these approaches often overlook scalability issues in real-time applications, where high-dimensional representations lead to computational bottlenecks [33].
The integration of deep learning has marked a paradigm shift in text-based content recommendation. A survey on deep learning-based recommender systems categorized models into neural collaborative filtering, autoencoders, and convolutional networks, showing their superiority in capturing non-linear relationships in text data [34]. More recently, foundation models have been adapted for recommendation, with a taxonomy covering feature-based, generative, and agentic paradigms [35]. These models excel in handling sequential user interactions, but their application to text-heavy content requires careful embedding strategies [36].
Transformer-based architectures have particularly revolutionized content recommendation by enabling better semantic understanding. The BERT-based Sequential Transformer (BST) model, for example, uses transformer layers to model user behavior sequences, achieving state-of-the-art performance on datasets like MovieLens [37]. Transformers4Rec, an open-source library bridging NLP and recommender systems, facilitates the use of pre-trained transformers for session-based recommendations [38]. In e-commerce, transformer models have been employed for substitute product recommendation, incorporating weakly supervised customer data to improve accuracy [39]. Music recommendation systems have also adopted transformers to capture sequential patterns in user listening histories [40]. A study on exploiting deep transformer models in review-based recommenders showed improved ratings prediction by processing textual reviews [41]. Furthermore, a transformer-based architecture for collaborative filtering, MetaBERTTransformer4Rec, outperformed baselines in personalized settings [42]. These works highlight the flexibility of transformers in text-based scenarios, but they often assume abundant resources, limiting applicability to diverse content types [47].
Scalability and efficiency remain critical challenges in transformer-based systems. Techniques like dimensionality reduction via Principal Component Analysis (PCA) or UMAP have been explored to compress embeddings for large-scale vector searches [22]. A survey on sequential recommendation systems discussed how transformers handle time-series data but emphasized the need for optimization in high-dimensional spaces [43]. In real-time applications, such as at Scribd, transformer models have been deployed for content recommendation, balancing accuracy with latency [44].
Despite these advancements, low-resource languages pose unique hurdles for text-based content recommendations Low-resource settings, characterized by limited labeled data and linguistic tools, hinder the development of effective NLP models [26]. A survey on Urdu language processing outlined challenges like morphological complexity and code-mixing, which affect semantic tasks [25]. In multilingual contexts, recommender systems often underperform for low-resource languages due to biased training data [45]. For instance, evaluations of large language models (LLMs) on low-resource languages like Bengali revealed gaps in zero-shot and few-shot prompting, relevant for recommendation personalization [46]. Multimodal approaches have been proposed to augment text-based systems in low-resource scenarios, but they require integrated frameworks [47].
Focusing on Urdu, a low-resource language with over 230 million speakers, text-based content recommendation remains underdeveloped. Existing research has primarily explored foundational techniques, often adapted from high-resource paradigms. Early systems relied on traditional lexical methods, such as TF-IDF with cosine similarity, which are limited by surface-level keyword overlap [14, 15]. A review of text-based recommendation systems discussed these challenges in low-resource languages like Urdu, advocating for advanced NLP to bridge semantic gaps [14]. Acknowledging this, subsequent work integrated pre-trained transformer models like BERT to generate semantic embeddings [6]. For Urdu content, a content-based model using UrduHack for preprocessing and BERT embeddings achieved higher accuracy in personalized recommendations [16].
Further innovations have enriched content representation through entity-based techniques. A foundational contribution in this direction was the development of an Urdu Wikification pipeline, which proposed a novel approach for linking named entities within Urdu text to their corresponding pages in Wikidata [17]. This work was pivotal for multiple reasons. First, it created and made publicly available valuable language-specific resources, including the SKEL dataset (550 annotated news titles) and the larger SKRS dataset (16,738 articles with user interaction data), which provided a critical infrastructure for subsequent research. Second, it demonstrated a practical methodology for constructing a sub-knowledge graph (8,439 entities and 23,080 relationships) from these linked entities. By applying the TransE algorithm to this graph to generate knowledge graph embeddings, the authors created a rich, structured semantic layer that could be integrated into a news recommendation model. The reported final accuracy of 60.8% for this hybrid RNN-based recommender established a new, knowledge-enhanced benchmark over purely lexical baselines for Urdu, showcasing the potential of leveraging structured external knowledge.
Building directly upon this wikification foundation, the later work SEEUNRS (Semantically Enriched Entity-Based Urdu News Recommendation System) refined and scaled the entity-centric approach [24]. This framework utilized an expanded dataset of 23,250 Urdu news articles to exploit the ”hidden semantic features” captured through entity and knowledge graph embeddings. The core advancement of SEEUNRS was its deeper integration of this entity knowledge to drive personalized recommendations. Rather than treating entities merely as tags, the system used their embeddings to model nuanced semantic relationships between users, articles, and concepts. This enabled the framework to outperform traditional lexical methods (e.g., TF-IDF) in precision, demonstrating that semantic enrichment through entities could more effectively address the vocabulary mismatch problem and capture user interest in low-resource settings.
Limitations of prior Urdu work include: (1) reliance on global, single-representation matching (no distinction between headline vs. full-text semantics); (2) no exploration of multiple pooling strategies or explicit evaluation of pooling choices; (3) absence of dimensionality-reduction and vector-storage engineering (no vector DB), which constrains large-scale retrieval and latency; and (4) no adaptive handling of query granularity (short vs. long queries are treated uniformly), hence not accommodating multiple use cases of recommendation results [27].
Our ULTRA framework addresses these gaps by introducing a dual-pathway system with query routing, optimized pooling, dimensionality reduction, and vector databases. This advances Urdu text-based content recommendation, generalizing beyond news to handle diverse textual corpora efficiently. Table I outlines the key limitations of prior work in Urdu text-based content recommendation systems and illustrates how our proposed ULTRA framework addresses them.
TABLE I: Comparative analysis of urdu text-based content recommendation systems and the proposed ULTRA framework
Ref. Year Model/Technique Dataset Paradigm Key Contributions Identified Limitations
Kanwal et al. [14] 2021 Literature Review / Survey N/A (Review Paper) Survey of Text-based RS Comprehensive review of methods; Identified core challenges for low-resource languages like Urdu. No implementation; highlights gaps but provides no solution.
Abbas et al. [16] 2022 TF-IDF + Cosine Similarity, BERT Embeddings Custom Urdu News (1,160 articles) Content-based Filtering Introduced semantic BERT embeddings for Urdu news, moving beyond pure lexical matching. Limited dataset; single-pipeline design; no adaptation for different query types; relies partly on lexical matching.
Kanwal et al. [17] 2022 Entity Linking + RNN Hybrid with TransE SKRS Dataset (16,738 articles) Knowledge Graph-based Hybrid Pioneered Urdu Wikification; used entity knowledge graph to enrich semantics for recommendation. Moderate accuracy (60.8%); complex pipeline; performance tied to entity presence; no query adaptation.
Kanwal et al. [24] 2024 Entity-Enriched Framework Expanded Dataset (23,250 articles) Entity-Based Semantic Recommendation Advanced entity-based matching, exploiting hidden semantic features for improved precision. Heavy reliance on entity recognition; no query-adaptive routing; limited scalability for real-time systems.
ULTRA 2026 Dual-Embedding Transformer with Query-Adaptive Routing Large-Scale Corpus (~112k articles) Adaptive Semantic Recommendation Novel dual-pathway architecture with length-threshold routing; optimized pooling & PCA; scalable vector DB integration. Addresses prior gaps: query-length variability, single-representation limitation, and scalability.
IV Proposed Methodology
In this research, we propose a methodology to develop a robust Urdu news recommendation system using semantic search powered by transformer-based embeddings. Figure 4 illustrates the architecture of the proposed system, and its modules are described below.
IV-A Problem Formulation
Given a user query in Urdu and a corpus of Urdu news articles, the objective is to retrieve the $k$ most semantically relevant articles. We formalize this as a semantic search problem in a high-dimensional vector space.
Let the article corpus be a set $D=\{d_{1},d_{2},\ldots,d_{m}\}$, where each article $d_{i}\in D$ is a tuple comprising its title and full textual content: $d_{i}=(h_{i},c_{i})$, with $h_{i}$ headline and $c_{i}$ the content/complete news artcile corresponding to that headline.
A user query is denoted by $q\in\mathcal{Q}$, where $\mathcal{Q}$ is the space of all possible Urdu text queries. The character length of a query is given by the function $\ell(q)$.
The core task of the system is to retrieve a set of articles that are semantically relevant to a user’s query. We model this using a binary relevance function $\text{rel}:\mathcal{Q}\times D\rightarrow\{0,1\}$. For a given query $q$ and article $d_{i}$, a human evaluator assigns:
$\text{rel}(q,d_{i})=\begin{cases}1,&\text{if article $d_{i}$ is relevant to query $q$,}\\ 0,&\text{otherwise.}\end{cases}$
A relevant article returned for a query constitutes a True Positive (TP).
Operationally, the system employs cosine similarity in a learned embedding space as an efficient proxy to approximate semantic relevance. It retrieves the top-$k$ articles, denoted $R_{q}^{k}$, ranked by this similarity score:
$R_{q}^{k}=\underset{d_{i}\in D,|R|=k}{\arg\max}\;\cos(\mathbf{v}^{\prime}_{q},\mathbf{v}^{\prime}_{i}).$ (1)
The primary performance metric is precision. For a single query $q$, the precision at $k$ (Precision@k) is the proportion of retrieved articles that are relevant:
$\text{Precision@k}(q)=\frac{1}{k}\sum_{d_{i}\in R_{q}^{k}}\text{rel}(q,d_{i})=\frac{|\text{TP}_{q}|}{|\text{TP}_{q}|+|\text{FP}_{q}|},$ (2)
where $\text{TP}_{q}$ is the set of True Positives and $\text{FP}_{q}$ is the set of False Positives (retrieved but irrelevant articles) for query $q$.
To evaluate the system’s overall effectiveness, we compute the mean precision across a test set of $N$ queries:
$\text{Mean Precision}=\frac{1}{N}\sum_{q\in Q_{\text{test}}}\text{Precision@k}(q).$ (3)
IV-A1 Embedding generation
Each textual element, be it headline or news article (named as content in our corpus), is mapped to a dense vector representation (embedding) using a pre-trained Urdu transformer model $\mathcal{M}$. The model generates a contextual embedding for each token. A pooling function $\mathcal{P}$ is applied to aggregate these token embeddings into a single, fixed-dimensional sentence embedding vector.
$\mathbf{v}_{q}$ $=\mathcal{P}(\mathcal{M}(q))\in\mathbb{R}^{768},$ (4)
$\mathbf{v}_{h_{i}}$ $=\mathcal{P}(\mathcal{M}(h_{i}))\in\mathbb{R}^{768},$ (5)
$\mathbf{v}_{c_{i}}$ $=\mathcal{P}(\mathcal{M}(c_{i}))\in\mathbb{R}^{768}.$ (6)
For the content $c_{i}$, if its token length exceeds the model’s maximum sequence limit $L_{\max}$, it is segmented into overlapping chunks $\{c_{i}^{(1)},\ldots,c_{i}^{(n)}\}$. The final content embedding is the average of the individual chunk embeddings:
$\mathbf{v}_{c_{i}}=\frac{1}{n}\sum_{j=1}^{n}\mathcal{P}(\mathcal{M}(c_{i}^{(j)})).$ (7)
IV-A2 Query-adaptive routing
To align the granularity of the query with the appropriate semantic representation, a threshold $\theta$ on query length governs the retrieval pathway:
$\text{Retrieval Space}=\begin{cases}\{\mathbf{v}_{h_{i}}\}_{i=1}^{m},&\ell(q)<\theta\\ \{\mathbf{v}_{c_{i}}\}_{i=1}^{m},&\ell(q)\geq\theta\end{cases}\ \ \begin{array}[]{l}\text{(Short Query Pathway)}\\ \text{(Long Query Pathway)}\end{array}$
The selection of the threshold $\theta$ is determined empirically based on the average length of titles or headings within the content corpus of the recommender system. In our case, corpus comprised of news articles, we approximated the headline length, given that headline embeddings ($\mathbf{v}_{h_{i}}$) are utilized for semantic comparison against short queries, and adopted this value as $\theta$ to delineate short from long queries. The underlying intuition for distinguishing between long and short queries and routing them to the appropriate pathway is already elaborated in Section II.
IV-A3 Dimensionality reduction for efficiency
To optimize large-scale retrieval, a dimensionality reduction function $\mathcal{R}:\mathbb{R}^{768}\rightarrow\mathbb{R}^{d^{\prime}}$, where $d^{\prime}\ll 768$, is applied to the pooled embeddings. For a given pathway, the reduced vectors are:
$\mathbf{v}^{\prime}_{*}=\mathcal{R}(\mathbf{v}_{*}),\quad*\in\{q,h_{i},c_{i}\}.$ (8)
This step preserves essential semantic relationships while significantly reducing the computational cost of similarity search.
IV-A4 Similarity computation and ranking
The semantic similarity between the query $q$ and an article $d_{i}$ is computed as the cosine similarity between their corresponding embeddings in the chosen (and potentially reduced) vector space. Let $\mathbf{v}^{\prime}_{q}$ be the query embedding and $\mathbf{v}^{\prime}_{i}$ be the target article embedding (either $\mathbf{v}^{\prime}_{h_{i}}$ or $\mathbf{v}^{\prime}_{c_{i}}$). The similarity score is:
$\text{sim}(q,d_{i})=\cos(\mathbf{v}^{\prime}_{q},\mathbf{v}^{\prime}_{i})=\frac{\mathbf{v}^{\prime}_{q}\cdot\mathbf{v}^{\prime}_{i}}{\|\mathbf{v}^{\prime}_{q}\|\|\mathbf{v}^{\prime}_{i}\|}.$ (9)
This cosine score serves as the operational retrieval metric. The final ranked list $R_{q}^{k}$ is obtained by sorting all articles in descending order by $\text{sim}(q,d_{i})$ and selecting the top $k$. It is important to note that while this geometric measure is an effective proxy for semantic relatedness, the final assessment of recommendation quality is based on human-evaluated relevance, as defined in Section IV-A. The system’s objective is to maximize the mean precision by optimizing the underlying embedding generation, routing, and dimensionality reduction components to produce $R_{q}^{k}$ that aligns as closely as possible with human relevance judgments.
This formulation provides a rigorous mathematical foundation for the subsequent development and evaluation of the ULTRA framework’s components, including the selection of $P$ (pooling strategies), the determination of $\theta$ (routing threshold), the choice of $R$ (dimensionality reduction technique), and the optimization of $d^{\prime}$ (reduced dimension). Notably, this problem formulation is general and can be applied to any recommendation system, although we specify news articles in our implementation due to the availability of the largest Urdu corpus dataset for evaluation. Nonetheless, the framework extends to a wide array of recommendation systems, encompassing digital libraries, news recommenders, product recommenders, and blog page recommenders.
IV-B Dataset
Our primary dataset was sourced from Kaggle [50]. We performed pre-processing on corpus of approximately 112k articles for our experiments. The dataset has 4 unique categories i.e., Science and Technology, Business and Economics, Entertainment and Sports. Figure 5 displays a high-level overview of the articles split across different categories.
Figure 5: Graphical representation of categories by frequency of Urdu news dataset
IV-C Dataset Preprocessing
The raw Urdu news data underwent a systematic pre-processing pipeline to ensure quality and consistency for subsequent analysis. This process consisted of five primary stages: (1) Dataset Loading and Encoding Normalization, where the initial corpus was loaded, and encoding issues were resolved to correct Urdu character representation. (2) Metadata Pruning and Composite Text Construction, involving the removal of non-informative attributes (e.g., author, date stamps not used in this study) and the concatenation of article headlines with corresponding body text to form a unified content field. (3) Text Cleaning and Normalization, which stripped HTML tags, punctuation, special characters, and extraneous whitespace to standardize the textual data. (4) Urdu Stop-Word Elimination, wherein a curated list of high-frequency, low-information Urdu words was filtered out to reduce noise. (5) Invalid and Redundant Record Filtering, where entries with null values, exact duplicates, or text length below a defined threshold were removed to finalize the dataset. Figure 6 illustrates provides a visual overview of the complete pre-processing pipeline.
Figure 6: Complete Urdu news dataset pre-processing pipeline
IV-C1 Dataset loading and encoding correction
Urdu text datasets often suffer from character encoding inconsistencies when processed through standard file reading methods. This issue arises due to variations in file encoding formats and platform-specific text handling. To preserve the integrity of the Nastaliq script, which includes complex character combinations and diacritical marks, the dataset was initially loaded using the ISO-8859-1 encoding scheme.
A custom decoding function was implemented to detect and correct misencoded characters programmatically. This function operates by detecting encoding mismatches through character pattern analysis, re-encoding problematic text segments from ISO-8859-1 to UTF-8, and validating the corrected text against Urdu Unicode ranges.
This systematic correction ensured accurate representation of all Urdu characters, including special diacritics (such as zabar, zer, and pesh) and punctuation marks specific to Urdu script. The encoding normalization was applied uniformly across all text-bearing columns: Headline, News Text, Category, and Source. This foundational step was critical to prevent information loss during subsequent processing stages.
IV-C2 Removal of irrelevant metadata and composite content creation
News datasets typically contain various metadata fields that, while potentially useful for some analysis, do not contribute to semantic understanding in text classification and embedding tasks. To reduce dimensionality and eliminate potential noise sources, non-informative metadata columns were systematically removed. These included: Index: Sequential identifiers with no semantic value, Date: Publication timestamps (not used in this text-focused study), URL: Web addresses containing platform-specific patterns, Source: Publication sources (retained only for provenance tracking), News Length: Derived statistics rather than raw content.
Following metadata pruning, a unified textual representation was created to capture comprehensive article semantics. The Headline and News Text columns were concatenated using a delimiter-aware joining method to form a new column named content. This composite field architecture preserves the concise, information-dense nature of headlines. It incorporates the detailed contextual information from article bodies and also creates a single input stream for embedding models, simplifying pipeline architecture.
The transformation ensures that models receive both the topic-signaling function of headlines and the explanatory depth of full articles. Figure 7 and Figure 8 illustrates the dataset structure before and after this transformation phase respectively.
Figure 7: Dataset before removal of irrelevant metadata
Figure 8: Dataset after removal of irrelevant metadata
IV-C3 Text cleaning and normalization
Raw text from web sources often contains structural artifacts, formatting remnants, and inconsistent character usage that can adversely affect downstream processing. To create a clean, standardized corpus, a multi-layered text normalization pipeline was implemented:
a)- HTML and formatting tag removal: Web-scraped content frequently includes HTML tags, CSS classes, and JavaScript fragments. These were stripped using a combination of regular expressions and HTML parsing libraries, ensuring complete removal while preserving textual content.
b)- Non-Urdu character elimination: Given the focus on Urdu-language analysis, non-relevant character sets were removed: English alphabets (A-Z, a-z), arabic numerals (0-9), mixed-language characters and symbols, platform-specific artifacts (e.g., emoticons, emojis not relevant to news context).
c)- Punctuation normalization: Both Urdu-specific and generic punctuation marks were systematically removed to create word-level tokens without boundary interference: generic punctuation: ‘.‘ (period), ‘,‘ (comma), ‘;‘ (semicolon), ‘:‘ (colon), quotation marks, brackets of all types, Urdu comma, Urdu semicolon, Urdu question mark.
d)- Whitespace standardization: Multiple consecutive whitespace characters (spaces, tabs, newlines) were collapsed into single spaces, and leading/trailing whitespace was trimmed. This normalization ensures consistent token boundaries during subsequent processing.
The combined effect of these cleaning operations is a homogeneous Urdu text corpus where semantic content is preserved while structural noise is minimized. This clean text serves as optimal input for both traditional NLP pipelines and modern embedding models.
IV-C4 Urdu stop-word elimination
Stop words (high-frequency function words that carry minimal semantic content) can dominate text representations without adding meaningful information. For Urdu, which has a rich set of grammatical particles and connectors, stop-word removal is particularly impactful. A manually curated list of 127 Urdu stop words was compiled through linguistic analysis and frequency distribution studies of the corpus. The stop-word list includes several categories of function words:
a)- Grammatical particles: Words that primarily serve grammatical functions rather than conveying content as shown in Figure 9.
Figure 9: Grammatical particles in Urdu language
b)- Conjunctions and connectors: Words that connect phrases but have limited standalone meaning. Figure 10 demonstrates some coordinating and subordinating conjunctions.
Figure 10: Conjunctions and connectors in Urdu language
c)- Demonstratives and pronouns: Common pronouns and demonstratives that appear frequently but lack specific content. Figure 11 gives some examples of demonstratives and pronouns.
Figure 11: Demonstratives and pronouns in Urdu language
d)- Common adverbs and discourse markers: Frequently occurring adverbs and discourse particles as in Figure 12.
Figure 12: Common adverbs and discourse markers in Urdu language
The removal process was applied after text cleaning but before any tokenization or feature extraction. This ordering ensures that stop words are identified in their normalized form, increasing removal accuracy. The efficacy of the Urdu stop-word removal process is demonstrated in Figure 13, which presents three comparative examples. Each case shows the preprocessed text (before), the resulting text after stop-word filtering (after), and the consequent reduction in text length (in characters), highlighting the substantial compression achieved while preserving core semantic content.
Figure 13: Common adverbs and discourse markers in Urdu language
Quantitatively, stop-word removal reduced the corpus size by approximately 33.5%, indicating the high frequency of these function words in Urdu news text. This reduction decreases computational requirements while potentially improving model performance by focusing on content-bearing terms. Figure 14 shows all 127 stop words that were identified and removed from dataset.
Figure 14: Common stop words in urdu language
IV-C5 Removal of invalid and redundant entries
Data quality assurance is critical for building reliable machine learning models. The preprocessing pipeline includes systematic filtering to eliminate problematic records through three validation criteria:
a)- Null value detection and removal: Records with missing values in essential columns were identified and removed. The validation focused on: content column: Empty or null text fields, Category column: Unlabeled or ambiguous category assignments, and Critical metadata: Missing values in fields required for analysis. This step ensures that all training examples contain complete information for both features and labels.
b)- Duplicate content identification: Duplicate articles, whether from republication, syndication, or crawling artifacts, can bias model training by over-representing certain content. This process maintains dataset diversity while eliminating redundant information.
c)- Minimum length filtering: Extremely short text entries (fewer than 15 characters) typically lack sufficient semantic content for meaningful analysis. These may include broken or truncated articles, navigational or interface text mistakenly captured, and placeholder or template content. Such entries were removed to ensure all training examples contain substantive content.
IV-C6 Post-Processing Dataset Structure
Following the complete pre-processing pipeline, the dataset is structured for optimal use in downstream tasks. The dataset is stored in a tabular format with UTF-8 encoding to ensure compatibility with Urdu text processing libraries and frameworks. This structured, cleaned dataset serves as the foundation for all subsequent tasks described in the following sections.
The combined effect of these filtering operations is a clean, consistent dataset optimized for machine learning applications. Table II summarizes the quantitative impact of the complete pre-processing pipeline on the dataset characteristics.
Metric Description After Preprocessing
Records Total number of articles 111,853
Article Length (avg.) Average characters per article 986.3
Headline Length (avg.) Average characters per headline 52.31
Words Total words processed 30,639,468
Stop Words High-frequency tokens removed 10,257,229
Removal Rate Percentage of stop words removed 33.5%
Categories Unique topic categories 5
Sources Distinct publication sources 6
Duplicates Redundant articles removed 8
TABLE II: Dataset statistics after preprocessing
IV-D Proposed Methodology
The ULTRA framework implements a semantic search-based recommendation system tailored for Urdu news. As formalized in Section IV-A, the core challenge is to retrieve semantically relevant articles from a corpus of over 112k Urdu news items based on user queries of varying lengths and semantic granularity. By ”query granularity,” we refer to the level of detail and specificity in a user’s search input—ranging from concise, keyword-like queries, to elaborate, descriptive sentences. However, in our framework, news articles or full-length documents predominantly function as long queries, targeted at the use case where recommendations are generated for content relevant to the article the user is currently viewing.
Our approach addresses this challenge through three interconnected components: (1) adaptive embedding generation that aligns with query granularity, (2) efficient dimensionality reduction for scalable retrieval, and (3) an optimized dual-path retrieval architecture. The complete methodology is designed to balance semantic accuracy with computational efficiency, making real-time recommendations feasible for large-scale deployment.
IV-E Adaptive Embedding Generation
The foundation of semantic retrieval lies in generating high-quality vector representations that capture the meaning of both queries and documents. We employ a pre-trained Urdu transformer model, urduhack/roberta-urdu-small, which outputs 768-dimensional contextualized embeddings for input tokens [51]. To obtain a fixed-dimensional representation for an entire text sequence, pooling strategies are required to aggregate token-level embeddings.
IV-E1 Pooling Strategy Analysis
Three common pooling techniques were implemented and evaluated: mean pooling, max pooling, and CLS pooling. Mean pooling computes the average of all token embeddings, effectively creating a centroid of semantic information. Max pooling extracts the maximum value across each dimension, potentially emphasizing salient features. CLS pooling utilizes the special classification token’s embedding, which is trained during pre-training to capture sentence-level semantics.
Initial evaluation revealed that the optimal pooling strategy depends on the nature of the text being encoded. For long-form article content, mean pooling demonstrated superior performance in capturing distributed semantic context. For concise headlines, CLS pooling proved most effective at preserving focused topical meaning. This observation informed the development of our dual-path embedding architecture.
IV-E2 Query-Adaptive Dual-Path Architecture
A key insight driving our methodology is the semantic granularity mismatch between user queries and article representations. Short, focused queries as in Figure 15.
Figure 15: Short length query example
typically express a specific topic or intent in few words, making them semantically more aligned with article headlines or titles of documents, which are themselves concise summaries, than with the distributed, detailed content of full articles. To address this, we implemented a query-adaptive routing mechanism based on character length, which dynamically selects the most appropriate semantic space for comparison.
Graphical analysis of headline length distribution of corpus in our case revealed a maximum of approximately 125 characters as illustrated in Figure 16.
Figure 16: Distribution of headline characters length in dataset
Therefore, to stay on safe side, we chose a slightly higher number, $\theta=150$ as the routing threshold. Queries with $\ell(q)<\theta$ are processed through the headline pathway, where the query embedding is compared against a pre-computed database of headline embeddings (that also stored full article and category as metadata) using cosine similarity. The top-$k$ most similar headlines are retrieved, and their corresponding full articles, stored as metadata, are recommended to the user. Queries with $\ell(q)\geq\theta$ are routed through the content pathway,where they are matched against embeddings of full article content (with headline and category as metadata in this case). The top-$k$ articles with the highest similarity scores are then retrieved and recommended to the user.
It is important to clarify that in the headline pathway, the system uses headline embeddings only for similarity matching. The actual recommendations presented to the user are the corresponding full articles. This design exploits the fact that headlines serve as accurate, dense semantic proxies for their articles, enabling efficient retrieval of relevant content without requiring comparisons against lengthy article bodies. The headline embedding for each article is stored alongside metadata containing its full content and category, ensuring that when a headline match occurs, the complete article is readily available for recommendation.
For content exceeding the model’s maximum sequence length of 512 tokens, a chunking strategy with 50-token overlap was implemented. The final content embedding was computed as the average of individual chunk embeddings, ensuring comprehensive semantic coverage of lengthy articles.
IV-F Dimensionality Reduction for Scalable Retrieval
With the corpus scaling to around 112,000 articles, performing nearest-neighbor search in the original 768-dimensional space becomes computationally intensive. To enhance scalability while preserving semantic relationships, we investigated four dimensionality reduction techniques: Principal Component Analysis (PCA), Uniform Manifold Approximation and Projection (UMAP), Linear Discriminant Analysis (LDA), and Autoencoders.
IV-F1 Evaluation Methodology
The performance of each reduction technique was assessed not in isolation, but by its fidelity to the original semantic space. For a given query, we compared the top-$k$ recommendations retrieved from the reduced-dimensional database against those from the original 768-dimensional database, which served as ground truth. The primary evaluation metric was the overlap score (the number of articles common to both result sets) providing a direct measure of how well the reduced space preserves retrieval rankings.
PCA, a linear technique maximizing variance preservation, consistently outperformed non-linear alternatives across both content and headline embeddings. Its deterministic nature and computational efficiency, combined with superior overlap scores, made it the preferred choice for both retrieval pathways.
IV-F2 Dimensionality Optimization
Following the selection of PCA, we optimized the reduced dimension $d^{\prime}$ to balance retrieval accuracy with computational efficiency. For content embeddings (long query pathway), we evaluated dimensions of 64, 128, and 256 across queries of varying lengths (150–350 characters). The 128-dimensional reduction provided the optimal trade-off, maintaining high overlap with ground-truth recommendations while significantly reducing query latency.
For headline embeddings (short query pathway), evaluation at 100 character length of 100 queries revealed that 64 dimensions sufficed to capture the semantic essence of concise headlines without noticeable degradation in retrieval quality. This dimension-specific optimization reflects the differing information density between lengthy articles and succinct headlines.
The pooled and reduced embeddings of optimal dimensions are then stored in two Chroma vector DBs: vector DB containing content column embeddings and vector DB containing headline column embeddings.
IV-G System Architecture and Implementation
The ULTRA framework implements the methodological components through an integrated pipeline illustrated in Figure 4. The system processes user queries through the following stages:
IV-G1 Query Processing and Routing
Upon receiving a user query $q$, the system first computes its character length $\ell(q)$. Based on the threshold $\theta=150$ (remember this $\theta$ is specific to our recommender system based on news corpus, and may vary across different recommender systems), the query is routed to the appropriate pathway. Both pathways employ the same base transformer model but differ in their text source, pooling strategy, and reduced dimension.
IV-G2 Embedding Generation and Retrieval
The system employs two separate vector databases, each tailored to a specific query type. The processing pipeline for a given query is determined solely by its character length $\ell(q)$ relative to the threshold $\theta=150$, which dictates the pooling strategy, dimensionality reduction configuration, and target database for similarity search.
For queries with $\ell(q)\geq\theta$ (long queries), the system follows the content pathway. The query text is encoded using mean pooling to generate a 768-dimensional embedding $\mathbf{v}_{q}$, which is then reduced to 128 dimensions via Principal Component Analysis (PCA). This reduced vector $\mathbf{v}^{\prime}_{q}$ is compared against a pre-computed database of content embeddings, each generated through the same mean pooling and PCA reduction process, using cosine similarity.
For queries with $\ell(q)<\theta$ (short queries), the system follows the headline pathway. The query undergoes CLS pooling to produce its initial embedding, which is then reduced to 64 dimensions via PCA. This compact representation is compared against a separate database of headline embeddings, each processed identically (CLS pooling + PCA to 64D), to find the most semantically similar headlines.
Both vector databases are implemented using ChromaDB with Hierarchical Navigable Small World (HNSW) indexing configured for cosine similarity, enabling efficient approximate nearest-neighbor search at scale.
IV-G3 Result Composition and Presentation
The top-$k$ articles ranked by cosine similarity are retrieved along with their metadata (headline, category, and for headline matches, the full news article). The final recommendations are presented to the user as a ranked list of news articles with relevant excerpts.
V Results and Discussion
This section presents a detailed analysis of the experimental results obtained from our Urdu news recommendation system. The analysis culminates in demonstrating the necessity of our proposed dual-embedding approach for handling the semantic disparity between short and long queries.
V-A Evaluation of Pooling Methods for Article Embeddings
To generate fixed-dimensional vector representations from the token-level embeddings produced by the model urduhack/roberta-urdu-small [51]., we implemented and compared three standard pooling strategies: mean pooling, max pooling, and CLS pooling. Each method aggregates information differently: mean pooling computes the average of all token embeddings in the sequence, capturing overall semantic content; max pooling selects the maximum value for each dimension across tokens, potentially emphasizing salient features; and CLS pooling utilizes the embedding of the special classification token ([CLS]), which is trained during pre-training to encapsulate sentence-level semantics.
Before proceeding to results section, it is highly recommended to clarify the distinction between short and long queries definition in our work in the light of use cases, as we have elaborated in Section II.
We used the pre-trained urduhack/roberta-urdu-small transformer model, which generates 768-dimensional token embeddings [51].. A fixed maximum sequence length of 512 tokens was employed. For long documents exceeding this limit, a chunking strategy with a 50-token overlap was applied, and the final embedding was computed as the average of all chunk embeddings. Text was tokenized with the model’s native tokenizer, applying standard padding and attention masking.
For each pooling method, we created separate vector databases containing embeddings of around 112k Urdu news articles, and stored each article’s headline and category as its metadata. To evaluate their retrieval effectiveness, we constructed a test set comprising 18 queries. However it is important to note that, in the last section, we presented a comprehensive evaluation of precision metrics across 100 test queries, thereby demonstrating the efficacy of our proposed methodology. In the current analysis, a smaller subset of queries was employed to delineate the performance disparities among various pooling methods, to facilitate the selection of the most effective one for subsequent implementation.
Queries were categorized into two distinct types based on length and semantic richness: 13 short queries (user typed search bar queries, concise, keyword-like phrases) as in Figure 17
Figure 17: Short length queries
and 5 long queries (could be user typed queries or complete articles, detailed, context-rich sentences or paragraphs) as demonstrated in Figure 18.
Figure 18: Long length queries
V-A1 Performance on Long Queries
For long, detailed queries, as in Figure 18, a different performance pattern emerged, as shown in Table III. Mean pooling achieved the highest average precision at 74%, followed by max pooling (50%) and CLS pooling (48%).
TABLE III: Precision@10 for long queries using different pooling methods on content embeddings
Query Mean Pooling Max Pooling CLS Pooling
TP Precision TP Precision TP Precision
Q1 9 0.9 (90%) 5 0.5 (50%) 4 0.4 (40%)
Q2 10 1.0 (100%) 10 1.0 (100%) 10 1.0 (100%)
Q3 4 0.4 (40%) 1 0.1 (10%) 0 0.0 (0%)
Q4 6 0.6 (60%) 5 0.5 (50%) 4 0.4 (40%)
Q5 8 0.8 (80%) 4 0.4 (40%) 6 0.6 (60%)
Average 0.74 0.50 0.48
Mean Precision 74% 50% 48%
The superiority of mean pooling for long queries can be attributed to its ability to aggregate semantic information from all tokens in the document. Long queries, which often contain multiple contextual clues and descriptive phrases, benefit from matching against a similarly comprehensive representation of article content. This distributed semantic matching proved more effective than the focused representations from CLS pooling or the potentially noisy salient features from max pooling. Figure 19 shows the visual analysis of different pooling methods for long queries. Interestingly, in later evaluations with an expanded test set of 100 queries and retrieval depth of $k=15$, mean pooling on content embeddings for long queries achieved precision scores exceeding 90%, confirming its robustness for detailed semantic matching.
Figure 19: Visual Analysis of precision@10 for long queries using different pooling methods on content embeddings
V-A2 Performance on Short Queries
Table IV presents the Precision@10 results for short queries shown in Figure 17 across the three pooling methods. CLS pooling demonstrated superior performance with an average precision of 60%, significantly outperforming mean pooling (53.07%) and max pooling (28.46%).
TABLE IV: Precision@10 for Short Queries using Different Pooling Methods on Content Embeddings
Query Mean Pooling Max Pooling CLS Pooling
TP Precision TP Precision TP Precision
Q1 8 0.8 1 0.1 10 1.0
Q2 0 0.0 1 0.1 3 0.3
Q3 10 1.0 10 1.0 10 1.0
Q4 3 0.3 0 0.0 5 0.5
Q5 1 0.1 3 0.3 4 0.4
Q6 7 0.7 10 1.0 3 0.3
Q7 8 0.8 3 0.3 7 0.7
Q8 1 0.1 0 0.0 4 0.4
Q9 9 0.9 0 0.0 9 0.9
Q10 6 0.6 0 0.0 6 0.6
Q11 10 1.0 9 0.9 10 1.0
Q12 4 0.4 0 0.0 6 0.6
Q13 2 0.2 0 0.0 1 0.1
Average 0.531 0.285 0.600
Mean Precision 53.07% 28.46% 60.00%
The performance variance highlights a critical insight: while CLS pooling achieved the highest overall precision for short queries, even its best performance (60%) was unsatisfactory for practical recommendation scenarios. Figure 19 shows the visual analysis of different pooling methods for short queries. This indicates a fundamental limitation of using content-based embeddings for short queries. The semantic representation of a comprehensive 500-word article is too coarse and diffuse to effectively match the focused intent encapsulated in a brief query.
Figure 20: Visual Analysis of precision@10 for short queries using different pooling methods on content embeddings
The persistent under-performance on short queries necessitated the innovative solution discussed in the following Section V-B, where we introduce headline embeddings specifically designed to bridge this semantic gap.
V-B Bridging the Semantic Gap: Headline Embeddings for Short Queries
The maximum precision of 60% for short queries, even with the best-performing pooling method on content embeddings, was deemed inadequate for a production-level recommendation system. This performance gap stems from a semantic granularity mismatch: short queries express focused user intent, while content embeddings represent the broad, multifaceted themes of full-length articles. This observation motivated the development of our dual-embedding approach, which seeks to align query semantics with appropriate document representations.
V-B1 Hypothesis and Implementation
We hypothesized that short user queries are semantically more similar to article headlines than to full article content. Headlines, by design, are concise, information-dense summaries that capture the core topic of an article, making them better semantic matches for similarly concise queries. To test this hypothesis, we generated a new set of embeddings from the headline column of our dataset using the same pre-trained model (urduhack/roberta-urdu-small [51].). For each pooling method (mean, max, and CLS), we created separate vector databases where each headline embedding was stored alongside its corresponding full article content and category as metadata. This design ensured that when a headline matched a short query semantically, the system could retrieve and recommend the complete article to the user, not just headline.
An important distinction in processing was that headlines, being inherently short, required no chunking and were embedded in a single forward pass of the transformer model, unlike lengthy article content which required the chunking strategy described earlier.
V-B2 Performance Evaluation on Short Queries
The evaluation focused exclusively on the same 13 short queries as illustrated in Figure 17 that were used for evaluation of pooling methods for content embeddings, as our intuition posited that the semantic compatibility between short queries and headlines would only be beneficial for this query type. For long queries, content embeddings remained the appropriate choice due to their comprehensive semantic coverage.
Table V presents the Precision@10 results for short queries shown in Figure 17 across the three pooling methods. CLS pooling demonstrated superior performance, significantly outperforming mean and max pooling.
TABLE V: Performance of Headline Embeddings with Different Pooling Methods on Short Queries (Precision@10)
Query Mean Pooling Max Pooling CLS Pooling
TP Prec. TP Prec. TP Prec.
Q1 8 0.8 4 0.4 9 0.9
Q2 6 0.6 3 0.3 9 0.9
Q3 7 0.7 10 1.0 10 1.0
Q4 1 0.1 0 0.0 8 0.8
Q5 6 0.6 6 0.6 9 0.9
Q6 7 0.7 8 0.8 9 0.9
Q7 7 0.7 5 0.5 10 1.0
Q8 2 0.2 0 0.0 9 0.9
Q9 9 0.9 9 0.9 9 0.9
Q10 4 0.4 5 0.5 8 0.8
Q11 9 0.9 6 0.6 9 0.9
Q12 3 0.3 3 0.3 9 0.9
Q13 7 0.7 4 0.4 6 0.6
Avg. TP 5.92 4.69 8.46
Avg. Prec. 0.5846 0.4846 0.8769
58.46% 48.46% 87.69%
Figure 21 shows the visual analysis of different pooling methods for short queries when embeddings of headline were generated.
Figure 21: Visual Analysis of precision@10 for short queries using different pooling methods on headline embeddings
Superiority of CLS Pooling on Headline Embeddings
For headline-based embeddings, CLS pooling achieved a remarkable average precision of 87.69%, significantly outperforming both mean pooling (58.46%) and max pooling (48.46%) on the same text source. This represents a 46.2% relative improvement over the best content-based method (CLS pooling at 60%). We posit that because headlines are short and syntactically focused, the [CLS] token, which is specifically trained during pre-training to capture sentence-level semantics, effectively encapsulates the entire headline’s meaning without the dilution that can occur from averaging across tokens (mean pooling) or the potential introduction of noise from selecting maximum values (max pooling).
V-B3 Case Analysis: Precision Breakthroughs
The headline embedding approach, particularly with CLS pooling, enabled breakthroughs on queries where other methods failed catastrophically. The single-word query in Figure 22 represents the most challenging case for semantic matching.
Figure 22: Urdu short query 8
Content embeddings struggled with this query, achieving only 0.4 precision with CLS pooling. In contrast, CLS-pooled headline embeddings achieved 0.9 precision, retrieving highly relevant technology-focused headlines such as shown in Figure 23, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 23: Urdu short query 8 recommendations
Similarly query 4 as in Figure 24 exemplifies how headline embeddings better capture domain-specific semantics.
Figure 24: Urdu short query 4
With content embeddings, precision was only 0.5, but headline embeddings with CLS pooling achieved 0.8 precision, correctly retrieving election-related headlines as shown in Figure 25, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 25: Urdu short query 4 recommendations
Query 5 as shown in Figure 26 demonstrates the improved specificity of headline embeddings.
Figure 26: Urdu short query 5
Previous methods retrieved general entertainment news from Bollywood and Hollywood due to the broad semantic representation in content embeddings. However, CLS-pooled headline embeddings correctly captured the nationality modifier ”Pakistani”, retrieving exclusively Pakistan-centric entertainment headlines achieving 0.9 precision as illustrated in Figure 27, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 27: Urdu short query 5 recommendations
Table VI provides a comprehensive analysis illustrating the dramatic improvements in precision for recommendation results generated in response to short queries, when employing CLS pooling on headline embeddings (which outperformed in this scenario) compared to CLS pooling on content embeddings (which excelled in the alternative case).
TABLE VI: Comparative Precision@10 for Short Queries: Content vs. Headline Embeddings
Query Content (CLS) Headline (CLS) Imp.
TP Prec. TP Prec.
Q1 10 1.00 9 0.90 $\mathbf{-10.0\%}$
Q2 3 0.30 9 0.90 $\mathbf{+200.0\%}$
Q3 10 1.00 10 1.00 $\mathbf{0.0\%}$
Q4 5 0.50 8 0.80 $\mathbf{+60.0\%}$
Q5 4 0.40 9 0.90 $\mathbf{+125.0\%}$
Q6 3 0.30 9 0.90 $\mathbf{+200.0\%}$
Q7 7 0.70 10 1.00 $\mathbf{+42.9\%}$
Q8 4 0.40 9 0.90 $\mathbf{+125.0\%}$
Q9 9 0.90 9 0.90 $\mathbf{0.0\%}$
Q10 6 0.60 8 0.80 $\mathbf{+33.3\%}$
Q11 10 1.00 9 0.90 $\mathbf{-10.0\%}$
Q12 6 0.60 9 0.90 $\mathbf{+50.0\%}$
Q13 1 0.10 6 0.60 $\mathbf{+500.0\%}$
Avg. Prec. 0.600 0.877
Overall Imp. +46.2%
Figure 28 shows the line chart comparison of all pooling methods using both content and headline embeddings.
Figure 28: Comparison of precision@10 for short queries of all pooling methods for both headline and content embeddings
V-B4 Implications for Pooling Strategy Selection
The experimental results yield clear guidelines for pooling strategy selection in our dual-embedding framework:
[…]
This optimized configuration leverages the strengths of each approach: mean pooling’s comprehensive aggregation for matching long, detailed queries against rich article content, yielding optimal results (74% precision) and CLS pooling’s focused semantic capture for matching short, intent-driven queries against concise headlines. The 87.69% precision achieved for short queries represents not just a quantitative improvement, but a qualitative transformation in the system’s ability to understand and respond to focused user intent.
The success of this approach validates our core hypothesis: aligning the semantic granularity of the query with an appropriately granular document representation is paramount for recommendation relevance. This insight forms the foundation of our proposed dual-embedding architecture.
V-C Dimensionality Reduction for Efficient Retrieval
While our dual-embedding approach established an optimal framework for semantic matching, operational efficiency remained a critical concern. With our dataset scaling to 112,000 records, performing nearest-neighbor searches in the original 768-dimensional space incurred significant computational and temporal costs. To enhance the scalability and real-time responsiveness of our recommendation system without substantially compromising retrieval quality, we investigated dimensionality reduction (DR) techniques. The objective was to find lower-dimensional representations that preserved the essential semantic relationships within both our content and headline embedding spaces.
V-C1 Dimensionality reduction for content embeddings
Given our established dual-embedding architecture as discussed in Section V-B4 , we first focused on the content embedding space, where mean-pooled 768-dimensional vectors represented full news articles. To identify the most suitable DR technique, we implemented and compared three prominent methods: Principal Component Analysis (PCA), Uniform Manifold Approximation and Projection (UMAP), and a neural Autoencoder.
Methodology and configurations
All techniques were configured to reduce embeddings from 768 dimensions to 64 dimensions, enabling a fair and direct comparison of their ability to preserve semantic information. The configurations were as follows:
  • PCA (Principal Component Analysis): A linear, unsupervised technique configured with n_components=64 to capture the directions of maximal variance in the original data.
  • UMAP (Uniform Manifold Approximation and Projection): A non-linear manifold learning method with hyperparameters n_neighbors=15, min_dist=0.1, and metric=’cosine’ to model the underlying topological structure.
  • Autoencoder: A symmetric neural network with encoder-decoder architecture (768 → 384 → 192 → 64 → 192 → 384 → 768). It was trained for 50 epochs using Mean Squared Error loss, the Adam optimizer (learning rate=0.001), employed Dropout (0.2) and Batch Normalization for regularization.
The selection of $K=50$ for evaluation was deliberate; while PCA and the Autoencoder showed comparable performance at smaller $K$ values (e.g., $K=10$), increasing the retrieval depth helped us expose which DR technique has superior ability to maintain ranking fidelity across a broader spectrum of similar items. The dimensionality reduction was applied to the 768-dimensional mean-pooled content embeddings, as mean pooling was established in Section  V-B4 as the optimal strategy for generating semantic representations of full-length articles.
Evaluation strategy
The performance of each DR method was evaluated by its fidelity to the original high-dimensional semantic space. For a set of 13 diverse, long-form test queries (representative of the ”long query” pathway), we retrieved the top-50 recommendations ($K=50$) from both the original 768-dimensional database (serving as ground truth) and each 64-dimensional reduced database. The key metric was the Overlap Score—the number of articles common to both result sets—expressed both as a raw count and a percentage of the retrieved items. We also computed the Jaccard Index ($J=|A\cap B|/|A\cup B|$) to measure the similarity between the recommendation sets. For a query $q$, with $A_{q}$ as the top-$k$ results from the original 768D space and $B_{q}$ as the top-$k$ results from the reduced dimensional space, the Jaccard Index at retrieval depth $k$ is defined as:
$J@k(q)=\frac{|A_{q}\cap B_{q}|}{|A_{q}\cup B_{q}|}=\frac{\text{Overlap}_{q}}{2k-\text{Overlap}_{q}}$ (11)
The overall effectiveness of a dimensionality reduction technique was then quantified by averaging this index across all $Q$ test queries, yielding the Mean Jaccard Index at $k$:
$\text{Mean-}J@k=\frac{1}{Q}\sum_{q=1}^{Q}J@k(q)$ (12)
This metric provides a normalized measure of set similarity that accounts for both true positives and the union of retrieved items, offering a robust assessment of how well the reduced space preserves the retrieval characteristics of the original embedding space.
Performance of DR techniques on content embeddings
Table VII presents the comparative performance of the three DR techniques on content embeddings. PCA demonstrated decisive superiority, achieving a mean overlap score of 74.92% and a mean Jaccard Index of 0.572.
TABLE VII: Performance of DR techniques on content embeddings (Top-50 Retrieval)
Query Overlap with 768D Ground Truth Overlap Count Jaccard Index
PCA UMAP AE PCA UMAP AE PCA UMAP AE
Q1 86.0% 46.0% 64.0% 43 23 32 0.754 0.299 0.471
Q2 86.0% 22.0% 36.0% 43 11 18 0.754 0.124 0.220
Q3 90.0% 46.0% 52.0% 45 23 26 0.818 0.299 0.351
Q4 62.0% 10.0% 16.0% 31 5 8 0.449 0.053 0.087
Q5 70.0% 46.0% 0.0% 35 23 0 0.538 0.299 0.000
Q6 42.0% 16.0% 18.0% 21 8 9 0.266 0.087 0.099
Q7 82.0% 32.0% 44.0% 41 16 22 0.695 0.190 0.282
Q8 78.0% 26.0% 44.0% 39 13 22 0.639 0.149 0.282
Q9 66.0% 28.0% 48.0% 33 14 24 0.493 0.163 0.316
Q10 92.0% 46.0% 78.0% 46 23 39 0.852 0.299 0.639
Q11 66.0% 16.0% 26.0% 33 8 13 0.493 0.087 0.149
Q12 76.0% 50.0% 74.0% 38 25 37 0.613 0.333 0.587
Q13 78.0% 38.0% 50.0% 39 19 25 0.639 0.235 0.333
Mean 74.92% 32.46% 42.31% 37.46 16.23 21.15 0.572 0.187 0.273
Figures 29 and 30 present the evaluation results of dimensionality reduction techniques applied to content embeddings evaluated on long test queries. Figure 29 shows a horizontal bar graph of the top-50 recommendation overlap between 64-dimensional reduced embeddings and the 768-dimensional ground truth, while Figure 30 displays the corresponding average Jaccard index distribution across techniques.
Figure 29: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth: overlap in top-50 recommendation across DR techniques for content vector DB
Figure 30: Comparative analysis of Jaccard Index distributions for evaluating DR technique fidelity to original semantic space
The results reveal a clear hierarchy: PCA significantly outperformed both UMAP and the Autoencoder. The linear projection of PCA proved exceptionally effective at preserving the semantic neighborhood structure crucial for ranking a large set of similar articles. While the Autoencoder showed reasonable performance on several queries, its inconsistency was notable (e.g., Q5 with 0% overlap), suggesting potential instability or overfitting during training. UMAP, despite its strength in visualizing complex manifolds, performed the poorest in this retrieval-focused task, indicating that its non-linear transformations may distort the global similarity relationships needed for effective nearest-neighbor search.
V-C2 Dimensionality reduction for headline embeddings
Following the conclusive results for content embeddings, we systematically applied the same evaluation framework to the headline embedding space, the critical component of our short query pathway. The identical three techniques (PCA, UMAP, and Autoencoder) were employed to reduce the 768-dimensional CLS-pooled headline embeddings to 64 dimensions, as CLS pooling was established in Section  V-B4 as the optimal strategy for generating semantic representations of headlines. The remaining configurations are same as were adopted in case of DR analysis for content embeddings, dicussed in Section  V-C1. The evaluation utilized the same 13 short queries as shown in Figure 17 that defined the short query category, with the original 768D headline embeddings serving as the ground truth. Retrieval depth was set to $K=50$ to provide a rigorous test of each technique’s ability to preserve ranking fidelity across a broad set of similar items.
V-C3 Results and Comparative Analysis
The performance metrics for headline embeddings are presented in Table VIII. PCA once again demonstrated clear superiority, achieving a mean overlap score of 65.10% and a mean Jaccard Index of 0.471.
TABLE VIII: Performance of DR techniques on headline embeddings (Top-50 Retrieval)
Query Overlap with 768D Ground Truth Overlap Count Jaccard Index
PCA UMAP AE PCA UMAP AE PCA UMAP AE
Q1 46.0% 10.0% 32.0% 23 5 16 0.299 0.053 0.190
Q2 20.0% 14.0% 18.0% 10 7 9 0.111 0.075 0.099
Q3 80.0% 30.0% 68.0% 40 15 34 0.667 0.176 0.515
Q4 62.0% 26.0% 40.0% 31 13 20 0.449 0.149 0.250
Q5 76.0% 20.0% 58.0% 38 10 29 0.613 0.111 0.408
Q6 82.0% 24.0% 50.0% 41 12 25 0.695 0.136 0.333
Q7 86.0% 12.0% 68.0% 43 6 34 0.754 0.064 0.515
Q8 66.0% 12.0% 58.0% 33 6 29 0.493 0.064 0.408
Q9 42.0% 14.0% 36.0% 21 7 18 0.266 0.075 0.220
Q10 66.0% 14.0% 50.0% 33 7 25 0.493 0.075 0.333
Q11 74.0% 18.0% 54.0% 37 9 27 0.587 0.099 0.370
Q12 84.0% 14.0% 62.0% 42 7 31 0.724 0.075 0.449
Q13 62.0% 20.0% 62.0% 31 10 31 0.449 0.111 0.449
Mean 65.10% 17.50% 50.46% 32.55 8.75 25.23 0.471 0.090 0.324
Figures 31 and 32 present the evaluation results of dimensionality reduction techniques applied to headline embeddings evaluated on short test queries defined in Figure 17 . Figure 31 shows a horizontal bar graph of the top-50 recommendation overlap between 64-dimensional ‘reduced embeddings and the 768-dimensional ground truth, while Figure 32 displays the corresponding average Jaccard index distribution across techniques.
Figure 31: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth for headline vector DB: overlap in top-50 across DR techniques
Figure 32: Comparative analysis of Jaccard Index distributions for evaluating DR technique fidelity to original semantic space for headline vector DB
V-C4 Selection of Dimensionality Reduction Technique
The Autoencoder showed improved relative performance on headline embeddings compared to content embeddings, yet remained inconsistent. UMAP continued to perform poorly, confirming its unsuitability for preserving global similarity structures in this retrieval context. PCA’s robust performance across two distinct textual domains (long-form content and short headlines) and two different optimal pooling strategies (mean and CLS) demonstrates its generalizability and reliability as a dimensionality reduction technique for semantic embedding spaces in information retrieval tasks.
Based on the comprehensive evaluation across both content and headline embeddings, Principal Component Analysis (PCA) was selected as the optimal dimensionality reduction technique for our dual-embedding architecture. The absolute overlap scores for headline embeddings (mean 65.10%) and content embeddings (mean 74.92%). The adoption of PCA enables efficient storage and rapid similarity search in our production system while preserving approximately 65-75% of the original semantic retrieval fidelity, as measured by overlap with full-dimensional results.
V-D Determining Optimal Embedding Dimension
Following the selection of Principal Component Analysis (PCA) as the superior dimensionality reduction technique, we sought to identify the optimal reduced dimension size for each pathway of our dual-embedding architecture. The primary objective was to maximize the preservation of semantic fidelity relative to the original 768-dimensional space while achieving a practical dimensionality for efficient storage and retrieval. This trade-off is inherent: lower dimensions may lead to greater information loss, while higher dimensions offer diminishing returns in semantic preservation at the cost of increased computational and storage overhead.
To systematically explore this trade-off, we evaluated three candidate dimensions, 64, 128, and 256 selected as powers of two for alignment with computational best practices in vector processing. The evaluation was conducted separately for content embeddings (serving long queries) and headline embeddings (serving short queries) to identify dimension-specific optima tailored to each semantic space.
V-D1 Optimal Dimension for Long Queries: Content Embeddings
In the preceding dimensionality reduction analysis, a fixed target dimension of 64D was used for all candidate techniques (PCA, UMAP, Autoencoder) to ensure a fair and consistent comparison of their ability to preserve semantic information. However, to optimize the system for practical deployment, it is essential to determine the specific reduced dimension that best balances retrieval speed with semantic fidelity for each embedding pathway. The trade-off is inherent: lower dimensions yield faster computations but risk greater information loss, while higher dimensions preserve more semantics at increased computational cost.
[…]
It is important to clarify the operational definition of a ”long query” in our system. Beyond traditional short search- bar queries, we consider the full text of an article a user is actively reading as a long query. This scenario is common in content recommendation systems, where the article being viewed serves as the contextual seed for generating related recommendations. In this case, the article’s entire content acts as the query, triggering the long-query pipeline. Empirical observation indicates that recommendation precision generally improves with query length, as richer context allows for more accurate semantic matching.
The lower bound for this pathway is $\theta$ = 150 characters. To ensure system robustness across the natural variation in article lengths, our evaluation for dimension selection was structured to test performance across a spectrum of long-query sizes. We constructed a test set of 100 queries at each of five distinct length intervals: 150, 200, 250, 300, and 350 characters, and repeated the evaluation for each interval to comprehensively assess performance across the anticipated range of input lengths. The mean-pooled content embeddings were compressed using PCA to each target dimension.
The key metric for evaluation was the Recommendation Overlap—the proportion of articles common between the top-$K$ results retrieved from the reduced-dimensional space and the top-$K$ results from the original 768D space.
Results and Analysis for Content Embeddings
The performance of PCA-reduced content embeddings across dimensions and query lengths is detailed in Table IX. The values represent averages computed across the 100 test queries for each length interval. Figures  33 and 34 shows grouped bars and heatmap analysis of statistics in table Table IX.
TABLE IX: Performance of PCA-reduced content embeddings across dimensions and query lengths
Query Length Average Overlap with 768D for 100 Test Queries(%)
(Characters) 64D 128D 256D
150 72.8 76.4 76.7
200 74.6 76.5 76.3
250 73.9 76.2 75.9
300 73.1 75.9 75.5
350 72.7 74.4 74.3
Mean 73.4 75.9 75.8
Figure 33: Bar graph analysis of overlap performance comparison across PCA dimensions and query lengths for content embeddings
Figure 34: Heatmap analysis of overlap performance comparison across PCA dimensions and query lengths for content embeddings
A consistent pattern emerged across all query-length scenarios: the 128-dimensional PCA-reduced embeddings achieved the highest mean overlap (75.9%), outperforming both 64D (73.4%) and 256D (75.8%). While 256D embeddings showed marginally higher overlap at the shortest query length (150 characters: 76.7% vs. 76.4%), this minimal advantage diminished with longer queries and reversed in the overall mean. The 64D embeddings consistently exhibited the lowest overlap across all intervals.
The results indicate that 128 dimensions provide an effective balance, capturing the essential semantic information of full-length articles with fidelity comparable to higher dimensions while avoiding the unnecessary complexity of 256D. Consequently, 128D was selected as the optimal reduced dimension for content-based retrieval within the long-query pathway.
V-D2 Optimal Dimension for Short Queries: Headline Embeddings
To determine the optimal reduced dimension for the short-query pathway, we applied a focused evaluation methodology to CLS-pooled headline embeddings. The evaluation considered the inherent characteristics of short queries, which are defined as having fewer than 150 characters and typically exhibit concise, focused semantic intent. Unlike long queries, which span a wide spectrum of lengths from the 150-character threshold to full article content, short queries demonstrate less variability in informational density and structure. Consequently, to determine an optimal dimension that generalizes across the short-query domain, we conducted the evaluation at a single, representative length of 100 characters. This length captures the typical upper bound of a concise search query or headline while remaining well within the short-query threshold, providing a robust benchmark for the compressed semantic space.
A test set of 100 distinct short queries was constructed at the 100-character length. The CLS-pooled headline embeddings were compressed using PCA to 64D, 128D, and 256D, and the performance of each was evaluated against the original 768-dimensional embeddings using the Recommendation Overlap metric, as defined previously.
Results and Analysis for Headline Embeddings
The results, presented in Table X, reveal a distinct optimum for the headline embedding space. The 64-dimensional PCA-reduced embeddings achieved the highest overlap score of 63.1%, surpassing both 128D (62.4%) and 256D (61.4%). Figure  35 shows bar graph and heatmap analysis of statistics in table Table X.
TABLE X: Performance of PCA-reduced headline embeddings across dimensions at 100-character query length
Query Length Average Overlap with 768D (%) Optimal Dimension
(Characters) 64D 128D 256D
100 63.1 62.4 61.4
Mean 63.1 62.4 61.4
Figure 35: Vertical bar chart showing average overlap percentages for different reduced dimensions (64D, 128D, 256D) for headline embeddings
This outcome indicates that for the semantically dense and focused space of headlines, a lower-dimensional representation (64D) is not only sufficient but optimal. The concise nature of headlines, where core meaning is concentrated in fewer tokens, allows for effective semantic capture in a more compact vector space. Increasing the dimension to 128D or 256D provides no improvement and, in fact, leads to marginally decreased overlap, likely due to the introduction of redundant or noisy dimensions that do not contribute meaningful variance for this specific data distribution.
Consequently, 64D was selected as the optimal reduced dimension for headline-based retrieval. This choice aligns with the principle of matching representation complexity to data complexity: the simpler, more focused semantics of headlines are best preserved in a lower-dimensional space, ensuring efficient storage and retrieval without compromising the fidelity required for accurate short-query recommendations.
V-E Final System Configuration and Performance Evaluation
Based on the comprehensive experimental analysis presented in above sections, the final retrieval pipeline of the ULTRA framework is configured as follows:
  • Content Retrieval Pathway (Long Queries, $\ell(q)\geq 150$ characters): Utilizes mean-pooled article embeddings reduced via PCA to 128 dimensions.
  • Headline Retrieval Pathway (Short Queries, $\ell(q)<150$ characters): Utilizes CLS-pooled headline embeddings reduced via PCA to 64 dimensions.
Table XI summarizes the complete configuration of this optimized dual-pathway architecture. The system dynamically routes queries to the appropriate semantic space based on their character length, with each space fine-tuned for maximum relevance through empirically validated pooling strategies and dimensionality reduction.
TABLE XI: Finalized Configuration of the Dual-Pathway Retrieval Pipeline
Component Content Path (Long Queries) Headline Path (Short Queries)
Source Text Article Content Article Headline
Pooling Strategy Mean Pooling CLS Pooling
Dimensionality Reduction PCA to 128 dimensions PCA to 64 dimensions
Vector Database ChromaDB (HNSW, cosine space) ChromaDB (HNSW, cosine space)
Key Metadata Headline, Category Full Content, Category
Optimal Use Case $\ell(q)\geq 150$ characters $\ell(q)<150$ characters
V-E1 Final Performance Assessment
To validate the effectiveness of the finalized system, we conducted a comprehensive performance assessment using a large-scale test set of 100 queries for each pathway. All true positives (relevant recommendations) were determined through manual evaluation by human assessors, following the rigorous methodology established in Section IV-A. The results, measured by Precision@15, demonstrate the system’s capability to deliver highly accurate recommendations.
Performance on Short Queries
For the headline retrieval pathway—designed for concise, keyword-focused queries typical of search bar inputs—the system achieved a mean Precision@15 of 94.35% for 100 queries. The distribution of precision scores across the 100 test queries was highly concentrated, with most queries achieving perfect precision (1.0). This performance represents a significant achievement, as short queries pose a fundamental challenge for semantic matching systems due to their limited contextual information. The 94.35% precision validates our core hypothesis: aligning short queries with semantically compatible headline embeddings effectively bridges the granularity gap that plagues traditional content-based approaches.
Performance on Long Queries
For the content retrieval pathway—serving detailed, context-rich queries or full article texts—the system achieved an exceptional mean Precision@15 of 98.53%. While this result is impressive, it aligns with expectations based on our experimental observations. Longer queries inherently contain richer semantic information, enabling more accurate embedding generation and subsequent similarity matching. The high precision achieved here validates the effectiveness of our optimized pipeline, combining mean pooling with PCA reduction to 128 dimensions, for preserving and matching comprehensive semantic contexts. The near-perfect performance across a diverse set of 100 test queries demonstrates the robustness and reliability of this pathway for applications such as ”read-next” article recommendations.
V-E2 Discussion of Results
The differential performance between pathways (94.35% for short queries vs. 98.53% for long queries) reflects the inherent difficulty of each recommendation task. The substantial improvement for short queries, from the initial 60% precision (in case of 13 test set queries) with content embeddings to 94.35% (in case of 100 test set queries) with headline embeddings, represents the primary contribution of our dual-embedding architecture. This transformation from unsatisfactory to production-grade performance addresses a critical limitation in semantic recommendation systems.
The exceptional precision for long queries, while expected given richer input context, nevertheless validates the optimization of each component in the pipeline: appropriate pooling, effective dimensionality reduction, and efficient vector search. The results across both pathways confirm that our systematic experimentation and component-wise optimization have yielded a robust, high-performance recommendation system capable of handling the full spectrum of user query types, from brief keyword searches to detailed contextual inputs.
VI Comparative Analysis with Baseline
Our experimental results demonstrate that the ULTRA framework achieves a Precision@15 of 0.9435 for short, keyword-based 100 queries routed through the headline embedding pathway. This performance is analyzed within the context of the existing landscape for Urdu content recommendation systems, which remains nascent.
VI-A Comparison with an Identified Baseline
A direct, albeit unpublished, baseline for Urdu news recommendation [16] employs a classical TF-IDF approach with cosine similarity. Their methodology involved preprocessing a corpus of approximately 1,160 articles categorized into Business, Science, Entertainment, and Sports. Their reported precision was 0.7143. A comparative summary is presented in Table XII.
TABLE XII: Comparison of the ULTRA Framework with a Baseline Urdu Recommender System
Aspect Baseline System [16] ULTRA Framework (This Work)
Core Methodology TF-IDF with Cosine Similarity Dual-Embedding Architecture (Headline CLS-pooling for $\ell(q)<\theta$; Content Mean-pooling for $\ell(q)\geq\theta$) + PCA
Reported Precision 0.7143 0.9435 (Precision@15 for short queries)
Evaluation Corpus Size $\sim$1,160 articles $\sim$118,000 articles
Relevance Judgment Not explicitly specified Human-evaluated True/False Positives for 100 test queries
Primary Contribution Applies classical IR and BERT embeddings to Urdu news Solves the semantic granularity mismatch; identifies optimal pooling and PCA-reduced dimensions (64D/128D)
The performance of our framework represents a substantial 32% relative improvement in precision over the identified baseline. This significant gain can be attributed directly to our core innovation: the dual-embedding architecture that resolves the semantic granularity mismatch by routing concise queries to a dedicated headline embedding space, optimized through empirical analysis of pooling strategies and PCA-based dimensionality reduction.
Furthermore, our evaluation was conducted on a corpus two orders of magnitude larger ($\sim$118k vs. $\sim$1.16k articles) and employed human-judged relevance assessments, providing a more robust and statistically reliable validation of the reported performance metrics.
VI-B Positioning within Urdu NLP Research
Our work addresses a distinct and relatively underexplored problem within the broader and active field of Urdu Natural Language Processing (NLP). Significant recent research has focused on and achieved high performance (>>92% accuracy/F1-score) in areas such as text classification for news categories, abstractive text summarization using transformer models like mT5, and large-scale fake news detection.
Prior analysis have noted that Urdu information retrieval systems have often relied on foundational techniques like TF-IDF and face challenges due to the language’s morphological complexity and resource scarcity. By successfully designing, optimizing, and validating a transformer-based semantic architecture specifically for content recommendation, this work provides a novel and high-performance pathway for advanced content discovery in Urdu, complementing the advancements in other NLP tasks.
VII Conclusions and Future Research Directions
This research presented and validated the ULTRA framework, a novel dual-embedding architecture designed to address the critical semantic granularity mismatch in Urdu news recommendation. Our core contribution is a system that dynamically routes queries based on character length: concise, keyword-driven queries ($\ell(q)<150$) are matched against CLS-pooled headline embeddings, while detailed, context-rich queries ($\ell(q)\geq 150$) are processed using mean-pooled content embeddings.
The experimental results conclusively demonstrate the efficacy of this approach. The most significant achievement is elevating the precision for short queries, a historically challenging task, from an unsatisfactory $\sim$60% using a single embedding space to a robust 94.35% (Precision@15) by aligning them with semantically compatible headline embeddings and evaluating for a larger test set queries. For long queries, the optimized pipeline achieved an exceptional 98.53% precision, validating the chosen combination of mean pooling and PCA-based dimensionality reduction to 128 dimensions. The systematic evaluation of pooling strategies and dimensionality reduction techniques established PCA as the optimal method for preserving semantic neighborhoods in both embedding spaces, with 64D and 128D identified as the respective optimal dimensions for headline and content pathways.
This work fills a notable gap in Urdu Natural Language Processing (NLP), where prior research has predominantly focused on foundational tasks like text classification and summarization. By developing and rigorously evaluating a sophisticated semantic recommendation system, we provide a validated, high-performance framework for Urdu content discovery.
The success of this framework opens several promising avenues for future work, both to address its limitations and to extend its capabilities. The most direct future direction is the deployment of this system as a production-grade application. A logical next step is to integrate the ULTRA framework into a dedicated Urdu news portal or aggregator. This application could feature a clean search interface for short queries and provide “Read Next” or “Related Articles” panels at the end of articles, powered by the long-query pathway. Moreover, the recommendation engine could be packaged as a cloud-based API service. This would allow existing Urdu digital publishers, blogging platforms, or digital libraries to integrate sophisticated semantic recommendation capabilities into their websites without developing the core NLP infrastructure in-house. The system can also be evolved into a hybrid recommender. The current semantic component could be combined with a collaborative filtering module that analyzes anonymized user interaction data. The final recommendation score could be a weighted combination of semantic similarity and personalized user relevance. Another interesting direction could be figuring out the same dual-embedding logic be applied to recommend related English articles for an Urdu query, or vice-versa, by using a multilingual embedding model. Furthermore, with the rise of multimedia news, could the system be extended to recommend relevant images or short video clips based on the semantic content of an article.
Acknowledgment
The authors acknowledge the use of the DeepSeek AI assistant as a supportive tool during the composition of this manuscript. It was utilized for sentence structuring, translation of ideas, and ensuring grammatical and formal English language generation, thereby aiding in the clear and effective communication of the research findings.